The perimeter of a rectangle is 24m and the area is 27sqm. Find its length and breadth.

Answer by ewatrrr(24785)

 
Hi,
.
P =2L + 2b = 24m
L + b = 12m
L = 12m-b
...
A = Lb
(12-b)b = 27m^2
b^2 -12b + 27 = 0
(b-9)(b-3)= 0
b = 3m 0r 9m and L = 9m 0r 3m
Dimensions are 3m by 9m